OpenStudy (anonymous):

How far does the tip of the minute hand of a clock move in 1 hour and 27 minutes if the hand is 2 inches long?

OpenStudy (callisto):

distance it moves \[=\pi(2)^2 + \pi(2)^2 ( 27/60)\] \[=4\pi + 1.8\pi \] \[=5.8\pi \] =18.22 inches